OpenStudy (anonymous):

The polynomial of degree 3, P(x), has a root of multiplicity 2 at x=2 and a root of multiplicity 1 at x=-2. The y-intercept is y= -4.8 . Find a formula for P(x) . Please check my answer! P(x)=(x-2)^2(x+2)-4.8 is that right?

OpenStudy (campbell_st):

nope not quite right \[P(x) = k( x+2)(x-2)^2\] then \[P(x) = k(x +2)(x^2 -4x + 4)\] so the constant term after distributing is 8 so you know that 4.8 = 8k solve for k then you can fully distribute to find the equation of the polynomial., Hope this helps.

OpenStudy (anonymous):

x^3-2x^2-4x-8 is 8 from this ? right?